Problem 9-7
Preferred stock rate of return

What will be the nominal rate of return on a perpetual preferred stock with a $100 par value, a stated dividend of 7% of par, and a current market price of (a) $61.00, (b) $85.00, (c) $95.00, and (d) $142.00? Round your answers to two decimal places.

a %

b%

c%

d %
Answer #1

Answer:

Calculation of Nominal rate of return :

Nominal Rate of return on a perpetual preferred stock = Annual Dividend / Market Price

a. Nominal Rate of return =(100*7%) / 61 = 0.1148 =11.48%

b. Nominal Rate of return =(100*7%) / 85 = 0.0824 =8.24%

c. Nominal Rate of return =(100*7%) / 95 = 0.0737 =7.37%

d. Nominal Rate of return =(100*7%) / 142 = 0.0430 =4.30%